OpenStudy (anonymous):

Jude has seven times as many dimes as nickels. Their value is $5.25. How many nickels does he have? 49 15 7

Parth (parthkohli):

Nickel = 5 cents Dime = 10 cents. Can you guess what equation would be used here?

OpenStudy (anonymous):

n+5=d+10

Parth (parthkohli):

But where is the 2.25 dollars thingy?

OpenStudy (anonymous):

i hate equations

Parth (parthkohli):

Convert 2.25 dollars to cents and you get 225. 225 will be on one side.

OpenStudy (anonymous):

u mean 5.25

Parth (parthkohli):

Let x be the number of nickels. And then just think how many dimes are there. 7 TIMES THE NUMBER OF DIMES. So 7x number of nickels.

Parth (parthkohli):

Yes I mean 5.25.

Parth (parthkohli):

Forming an equation, you'll get: \( \color{Black}{\Rightarrow x(10) + 7x(5) = 525 }\) x times 10 cents plus 7x times 5 cents will make 525. Simple as that.

Parth (parthkohli):

Do you know how to solve this equation?

OpenStudy (anonymous):

ill try ^-^

Parth (parthkohli):

First step would be to distribute by the way.

Parth (parthkohli):

What is 10 times x? What is 7x times 5?

OpenStudy (anonymous):

10x and 35x

OpenStudy (anonymous):

=45x

Parth (parthkohli):

Sorry, the equation would be: \( \color{Black}{\Rightarrow 7x(10) + 5x = 525}\)

OpenStudy (anonymous):

70x+5x=525

OpenStudy (anonymous):

=75x

Parth (parthkohli):

Yes, add 70x and 5x. 75x = 525. What will be the next step?

OpenStudy (anonymous):

i dont know my brain hurts

Parth (parthkohli):

To isolate the variable x, you may divide both sides by 75.

OpenStudy (anonymous):

x=7

Parth (parthkohli):

To isolate any variable, do the opposite operation already done. \( \color{Black}{\Rightarrow x + 3 = 90 }\) For example ^. You'll subtract both sides by 3.

Parth (parthkohli):

Yes, and there are x nickels.

OpenStudy (anonymous):

oh yaa its coming back to me now

OpenStudy (anonymous):

that part anyway

Parth (parthkohli):

x = 7, so 7 nickels.

OpenStudy (anonymous):

yup

Parth (parthkohli):

You got it?

OpenStudy (anonymous):

yup